OpenStudy (cherylim23):

For the function f(x,y) = [sin(x-y)]^2, find the maximum and minimum rate of change and in which direction it occurs. Also find the direction for which Du f is zero. I've got the gradient function from there, which is 2sin(x-y)cos(x-y) i - 2sin(x-y)cos(x-y) j but no sure how to progress from there...

OpenStudy (anonymous):

\[\left| \Delta f \right|=\sqrt{2} \ \left| \sin(2x-2y) \right|\] Delta is gradiant

OpenStudy (anonymous):

maximum of gradient is for when \[\sin (2x-2y)=\pm 1\] and the minimum of it for \[\sin (2x-2y)=0\]

OpenStudy (cherylim23):

Thanks, how do you find the direction in which it occurred? (i.e. the vector?)

OpenStudy (anonymous):

i think depends on (x,y) for example maximum of it occurs in direction \[\pm(i-j) \\ since \\ \sin(2x-2y)=\pm1\]

OpenStudy (anonymous):

can u imagine the places for which the minimums take place look likes to what?

OpenStudy (cherylim23):

Like a local (more like, global) minimum point as in a hyperbolloid?

OpenStudy (anonymous):